OpenStudy (anonymous):

given that Earth's surface gravitational field strength has a magnitude of 9.8N/kg, determine the distance above Earth's surface at which the magnitude of the field strength is 3.2N/kg

OpenStudy (anonymous):

\[F=Gm _{1}m _{2}/r ^{2}\] use this to see that the force changes invers proportionaly to square of distance

OpenStudy (anonymous):

m1 is mass of earth? G is gravitational constant? what is m2?

OpenStudy (anonymous):

m2 would be one kg, but you don't need that. 9,8/r2 = 3.2

OpenStudy (anonymous):

the answer is 0.75.

OpenStudy (anonymous):

sqroot(9.8/3.2)=1.75

OpenStudy (anonymous):

i would say 1,75 earth radius

OpenStudy (anonymous):

M= earth mass m= 1 kg G = universal constant r = distance \[9,8 = GMm/r _{1}^{2}\] \[3.2 = GMm/r _{2}^{2}\] deviding bouth: \[9,8/3,2 = r _{2}^{2}/ r _{1}^{1}\rightarrow 1,75 r _{1} =r _{2}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

give medal ?